When 300.0 g of a molecular compound was dissolved in 2000. g of water, the freezing point of the solution was lowered to -3.00 °C. What is the molecular weight of the compound?

 



1
Expert's answer
2021-02-03T07:24:38-0500

Freezing point of solution(T)= -3.00 °c

Freezing point of pure water(T°) = 0.00 °C

Mass of water = 2000 g = 2000/1000 = 2.0 kg

Mass of solute = 300.0 g

Moles of solute = 300.00/MM

Where, MM = Molecular weight

Now using the formula ,

T° - T = Kb × m

Where, m = Molality= mole/mass(Kg), Kb = freezing constant = 1.86 °c/m

0.0 - (-3.0) = 1.86 × 300/MM × 2.0

MM= 93.0 g/mol

So, molecular weight = [93.0] g/mol